OpenStudy (i_love_my_nieces):

You borrow $1,200 at a rate of 4.5% for a period of 6 months. How much will you repay at the end of the 6 months? 1. $27 2. $54 3. $1,173 4. $1,227

OpenStudy (i_love_my_nieces):

I believe it's $54.

OpenStudy (i_love_my_nieces):

@luisdabroski

OpenStudy (i_love_my_nieces):

@robert200019

OpenStudy (i_love_my_nieces):

@secret66

OpenStudy (inkyvoyd):

I'm pretty sure you have o repay more than you borrow :)

OpenStudy (inkyvoyd):

otherwise we would borrow money a lot more often :)

OpenStudy (i_love_my_nieces):

@inkyvoyd can you help me then?

OpenStudy (inkyvoyd):

hmm. you borrowed 1200. you have to return more than 1200. which of those numbers makes sense then?

OpenStudy (i_love_my_nieces):

D?

OpenStudy (inkyvoyd):

yeah, I would say.

OpenStudy (anonymous):

amount=principal (1+4.5/200)^1=1200(204.5/200) =1200*2045/2000 =12*2045/20=1227.00rs